OpenStudy (anonymous):

(x^2 - 4) (x - 1) anybody?

OpenStudy (anonymous):

What does x equal?

OpenStudy (anonymous):

@lacedupkid

OpenStudy (campbell_st):

you can split the 1st factor and rewrite it as \[(x^2 - 4)(x -1) = x^2 \times (x -1) - 4 \times (x - 1)\] then distribute.